Trägheitsmomente in Bezug auf Springer: HOHLZYLINDER: J_{z} = \frac{1}{2} m(r_{a}^{2} + r_{i}^{2}), J_{u} = \frac{m}{4} (r_{a}^{2} + r_{i}^{2} + \frac{h2}{3} ) HOHLKUGEL: J = \frac{2}{5} m \frac{r5 a−r5 i}{r3 a−r3 i} KREISBOGEN: Wie Sektor mal 2 TORUS: J = m(r_{T orus}^{2} + \frac{3}{4} r_{Dicke des Torus}^{2}) bei Rot. um Achse durch CM, ⊥ auf Torusebene